OpenStudy (anonymous):

The sum of the first 17 terms of an arithmetic sequence is 1,360. If the first term is 16, find the 17th term.

OpenStudy (amistre64):

ohh.... SUM of the first terms .... i had it a little offish

OpenStudy (anonymous):

use the formula Sn = n/2 [ a + l ]

OpenStudy (anonymous):

What does everything stand for?

OpenStudy (anonymous):

ok Sn = sum of n terms , n = number of terms, a = first term and l = last term. you have value of everything except l which is what you need to find. n=17, a = 17 and Sn = 1360

OpenStudy (anonymous):

sorry a = 16

OpenStudy (anonymous):

r u happy with solving the formula for l ?

OpenStudy (anonymous):

I'm having a hard time :/

OpenStudy (amistre64):

the sum of an arthritic series is rather ingeniously: number of terms (first term + last term) --------------------------------- = sum of series 2 i believe

OpenStudy (amistre64):

for example: 3+6+9+12 should equal: 4 (3+12) ------- = 2(15) = 30 2 18 12 --- 30

OpenStudy (anonymous):

I answered it in the previous post..:-)

OpenStudy (amistre64):

so we area given: 17 terms of an arithmetic sequence is 1,360. If the first term is 16, find the 17th term. 1360 = 17 (16 + A)/2 1360(2) = 17 (16 + A) 1360(2) ------- = 16 + A 17 1360(2) ------- - 16 + A 17

OpenStudy (amistre64):

1360(2) ------- - 16 = A ; where A is the last term, in this case #17 17

OpenStudy (amistre64):

.... if you come to 144 youll do good ;)

OpenStudy (anonymous):

Or the next post....:-)

OpenStudy (anonymous):

Omg I didn't check back on that estudier. Thank you guys so much :)